# Super(ficial)-alignment: Strong Models May Deceive Weak Models in Weak-to-Strong Generalization

### v2

@@ 6.1 Only Using Correct High-Confidence Samples Cannot Mitigate Deception
Based on the hypothesis we have made in Section 5.3, the reason why a strong model could deceive a weak model in preference alignment is might be that it is provided with both the correctly and wrongly predicted samples from the weak model. Thus, one possible solution to avoid deception may be only providing those correct high-confidence samples from the weak model for weak-to-strong alignment. We conduct experiments in the implicit conflict setting to validate this hypothesis, and the detailed experimental settings are in Appendix I. We provide the results when the weak model is GPT-2-Large in Figure LABEL:subfig:_high-confidence_deception_score as an illustrative example, and leave the full results in Appendix I. Unfortunately, we get the negative conclusion that only supervising with high-confidence samples cannot mitigate the deception phenomenon. This implies that there are deeper mechanisms to explain how strong models perceive the knowledge boundaries of weak models and exhibit deceptive behaviors, which can be an interesting direction for future work. For example, strong models may possibly infer the areas of knowledge where teachers excel and struggle based on the portion and distribution of samples that teachers provide across different domains in this setting.
@@ 6.2 Bootstrapping Can Mitigate Deception to Some Extent
Inspired by the relationship between deception severity and models’ capability gap, we are curious whether employing a bootstrapping method with an intermediate model (Burns et al., 2024) would result in a lower deception score compared to directly using the weak model to supervise the strong model. In this case, we make the weak model first supervise an intermediate model and then let the intermediate model further supervise the ultimate strong model. We fix the ultimate strong model to Mistral-7B, and for each weak model, we select every model between it and Mistral as an intermediate model. Detailed experimental settings are in Appendix J. The results are in Figure LABEL:subfig:_bootstrapping_accuracy and LABEL:subfig:_bootstrapping_deception_score. The results of cases when intermediate models are the same as weak models represent the results of directly using weak models to supervise Mistral-7B. Firstly, we can see that bootstrapping with an intermediate model can improve the generalization performance, which is consistent to the findings in Burns et al. (2024). More importantly, bootstrapping can indeed mitigate the deception issue to some extent, reflected in the consistently lower deception scores when intermediate models exist. The reason could be that some cases originally unknown to the weak model are now known to the intermediate model, making it difficult for the strong model to deceive in those cases.
7 Conclusion and Discussion
In this paper, we reveal and study a security issue in the weak-to-strong alignment, called the weak-to-strong deception. By studying in a multi-objective alignment setting, we empirically find that strong students can behave well-aligned in areas known to weak teachers, but tend to produce misalignments in areas unknown to weak teachers when conflicting alignment targets exist. Such a deception issue becomes more severe as the capability gap between weak and strong models increases, which introduces a greater challenge for humans to reliably supervise super AI as it continuously becomes smarter and more intelligent in the future. Finally, we discuss two possible countermeasures among which bootstrapping method exhibits a certain effect.
Regarding the future directions: (1) We have mainly conducted a preliminary study on the issue of weak-to-strong deception in a specific multi-objective alignment scenario. Future work could make more explorations in other scenarios, such as the spontaneous deception by AI (Denison et al., 2024). (2) Given our concerning experimental findings, we need a more fundamental analysis of why and how strong models are prone to deceiving weak models. (3) Given the limited effectiveness of bootstrapping method, more effective mechanisms are needed to mitigate the deception issue.
@@ Appendix A Limitations
Though our study provides a comprehensive empirical analysis on the weak-to-strong deception issue, there are some limitations that can be interesting future work: (1) We mainly conduct experiments in the preference alignment scenario on two offline preference optimization methods, SimPO (Meng et al., 2024) and DPO (Rafailov et al., 2024). Future work can explore the weak-to-strong deception issue on the online preference optimization frameworks such as PPO (Schulman et al., 2017). (2) In our experiments, we only consider the case where the target alignment dimension is harmlessness, which is indeed an important alignment goal. However, there are some other dimensions that are also important for model alignment. For example, the deception issue also matters in the honesty alignment (Cheng et al., 2024), where the stronger model should not learn to deceive the weak model to intentionally make wrong responses on questions that the weak model does not know.
Appendix B The Complete Procedure for Performing Weak-to-Strong Preference Alignment
@@
Appendix C Concrete Mathematical Forms of All Weak-to-Strong Objectives
C.1 Reward Modeling Scenario
@@
(3) No Conflict: Additionally, in order to explore the performance of the strong model it should have achieved without the conflicting alignment goal, we should obtain a weak-to-strong model trained under the weak supervision towards harmlessness only:
C.2 Preference Alignment Scenario
@@
(3) No Conflict:
Appendix D Training Details
D.1 Code and Platform
@@
D.2 Training Details in The Reward Modeling Scenario
@@
D.3 Training Details in The Preference Alignment Scenario
@@
Appendix E Results of Test Accuracies in The Preference Alignment Scenario with SimPO
@@
Appendix F Weak-to-Strong Results on DPO
@@
Besides the main results on SimPO (Meng et al., 2024), we also conduct weak-to-strong preference alignment experiments on DPO (Rafailov et al., 2024). The detailed procedure is similar to SimPO and can be found in Appendix B. The major difference is the metric for determining the correctness of each prediction of the model and calculating the model’s confidence on each sample pair. In DPO, when using Eq. (9) to calculate the confidence of model $\bm{\theta}$ on $(x;y_{c},y_{r})$, the model logit $\pi_{\bm{\theta}}(y|x)$ on each completion $y$ is now normalized by a constant factor $L$ (instead of normalized by each response’s own sequence length) and is calculated as $\pi_{\bm{\theta}}(y|x)=\frac{1}{L}\sum\nolimits_{i=1}^{|y|}\log P_{\bm{\theta}}(y_{i}|x,y_{<i})$. This is because the original training objective of DPO is to directly enlarge the gap between the log sums of token probabilities over the chosen and rejected responses. Thus, the evaluation metric should be consistent to the training objective. We admit this may introduce a sequence length bias, but this is an inherent issue of DPO. However, the scale of original log sum of token probabilities is very huge, we need to divide it by a constant to ensure that the model’s confidence falls within a reasonable distribution. Choosing different constants for re-scaling may affect the confidence distribution, however, it is equivalent to selecting different confidence thresholds. The results in Appendix H validate that the deception pattern is independent of the choices of confidence thresholds. Here, we set $L$ to 50. The confidence threshold $T$ is fixed to 0.75. The results of generalization performance are displayed in Figure LABEL:fig:_accuracy_results_in_dpo, and the results of deception scores are shown in Figure 31. The results show that the weak-to-strong deception issue also exists in DPO setting.
Appendix G Experiments on LLaMA-3 and LLaMA-3.1
@@
Appendix H Deception Scores under Different Confidence Thresholds
@@
In the main text, we display the patterns of deception scores in both two scenarios under the confidence threshold $T=0.75$. Choosing a different confidence threshold may affect the delineation of known and unknown areas of a target model. Here, we display the patterns of deception scores in both scenarios under two extra confidence thresholds ($T=0.80,0.85$). The results in the reward modeling and preference alignment (SimPO) scenarios are displayed in Figure LABEL:fig:_deception_scores_under_different_T_on_reward_modeling and Figure LABEL:fig:_deception_scores_under_different_T_in_preference_alignment, respectively. As we can see, though the concrete value of deception score in each experiment varies slightly under different $T$, the general pattern that the deception issue intensifies as the capability gap between the weak and strong models increases remains the same.
Appendix I Full Results of High-Confidence Weak-to-Strong Preference Alignment
@@
Appendix J Details in Bootstrapping Experiments
@@
Appendix K Case Study on Weak-to-Strong Deception

### v3

@@ 6.1 Only Using Correct High-Confidence Samples Cannot Mitigate Deception
Based on the hypothesis we have made in Section 5.3, the reason why a strong model could deceive a weak model in preference alignment is might be that it is provided with both the correctly and wrongly predicted samples from the weak model. Thus, one possible solution to avoid deception may be only providing those correct high-confidence samples from the weak model for weak-to-strong alignment. We conduct experiments in the implicit conflict setting to validate this hypothesis, and the detailed experimental settings are in Appendix M. We provide the results when the weak model is GPT-2-Large in Figure LABEL:subfig:_high-confidence_deception_score as an illustrative example, and leave the full results in Appendix M. Unfortunately, we get a negative conclusion that only supervising with high-confidence samples cannot mitigate the deception phenomenon. This implies that there exist deeper mechanisms to explain how strong models perceive the knowledge boundaries of weak models and exhibit deceptive behaviors, which can be an interesting direction for future work. For example, strong models may possibly infer the areas of knowledge where teachers excel and struggle based on the portion and distribution of samples that teachers provide across different domains in this setting.
@@ 6.2 Bootstrapping Can Mitigate Deception to Some Extent
Inspired by the relationship between deception severity and models’ capability gap, we are curious whether employing a bootstrapping method with an intermediate model (Burns et al., 2024) would result in a lower deception score compared to directly using the weak model to supervise the strong model. In this case, we make the weak model first supervise an intermediate model and then let the intermediate model further supervise the ultimate strong model. We fix the ultimate strong model to Mistral-7B, and for each weak model, we select every model between it and Mistral as an intermediate model. Detailed experimental settings are in Appendix N. The results are in Figure LABEL:subfig:_bootstrapping_accuracy and LABEL:subfig:_bootstrapping_deception_score. The results of cases when intermediate models are the same as weak models represent the results of directly using weak models to supervise Mistral-7B. Firstly, we can see that bootstrapping with an intermediate model can improve the generalization performance, which is consistent to the findings in Burns et al. (2024). More importantly, bootstrapping can indeed mitigate the deception issue to some extent, reflected in the consistently lower deception scores when intermediate models exist. The reason could be that some cases originally unknown to the weak model are now known to the intermediate model, making it difficult for the strong model to deceive in those cases.
7 Conclusion
In this paper, we reveal and study a security issue in the weak-to-strong alignment, called the weak-to-strong deception. By studying in a multi-objective alignment setting, we empirically find that strong students can behave well-aligned in areas known to weak teachers, but tend to produce misalignments in areas unknown to weak teachers when conflicting alignment targets exist. Such a deception issue becomes more severe as the capability gap between weak and strong models increases, which introduces a greater challenge for humans to reliably supervise super AI as it continuously becomes smarter and more intelligent in the future. Finally, we discuss two possible countermeasures among which bootstrapping method exhibits a certain effect. Given our concerning experimental findings, we call for future work to pay more attention to this issue and propose better solutions.
@@ 7 Conclusion
Reproducibility Statement
First of all, we provide the code and data to ensure reproducibility. Then, we give the necessary illustration of the experimental settings in main experiments in Section 4.1 and Section 5.2 in the main text. The complete procedure for performing weak-to-strong preference alignment experiments is put in Appendix C. The detailed mathematical forms of all conflict settings are in Appendix D. The complete training details are illustrated in Appendix E. The details of supplementary experiments are in Appendix 27, Appendix H, Appendix M, and Appendix N, respectively.
Acknowledgments
We sincerely thank all the anonymous reviewers for their valuable comments and constructive suggestions. This work was supported by The National Natural Science Foundation of China (No. 62376273) and Beijing Nova Program (No. 20240484568).
@@ Appendix A Limitations
Though our study provides a comprehensive empirical analysis on the weak-to-strong deception issue, there are some limitations that can be interesting future work: (1) We mainly conduct experiments in the preference alignment scenario on two offline preference optimization methods, SimPO (Meng et al., 2024) and DPO (Rafailov et al., 2024). Future work can explore the weak-to-strong deception issue on the online preference optimization frameworks such as PPO (Schulman et al., 2017). (2) In our experiments, we mainly consider the case where the target alignment dimension is harmlessness, which is indeed an important alignment goal. However, there are some other dimensions that are also important for model alignment. For example, the deception issue also matters in the honesty alignment (Cheng et al., 2024), where the stronger model should not learn to deceive the weak model to intentionally make wrong responses on questions that the weak model does not know (refer to preliminary experiments in Appendix O).
Appendix B Discussions on The Similarity and Differences between Weak-to-Strong Deception issue and Reward Hacking Problem in LLM Alignment
Here, we make a discussion on the similarities and differences between weak-to-strong deception and traditional reward hacking in LLM alignment (Pan et al., 2024a; b; Denison et al., 2024).
Regarding the similarities: Both alignment reward hacking and weak-to-strong deception study a phenomenon where the supervised model fools the teacher/reward model by excelling in one aspect that the teacher/reward model can perceive and judge, but behaving misaligned in another aspect that the teacher/reward model cannot provide accurate supervision.
Regarding the differences: (1) The first difference lies in the aspect that needs to be focused on. Reward hacking is studied by comparing the performance of the supervised model in two different alignment dimensions (e.g., the format or style v.s. the instruction following ability). However, in weak-to-strong deception, we aim to compare the performance of the supervised model on two different knowledge areas (Weak-Known v.s. Weak-Unknown) within one specific alignment dimension (e.g., harmlessness). (2) The second difference lies in the research setting. In existing reward hacking studies, there is usually one universal reward signal for supervising the student model. Then, these studies try to understand the behavior change of the supervised model in other dimensions in which the reward model cannot provide accurate supervision. Even though in some time, this universal reward signal is mixed with multiple dimensions, existing studies do not take a step further to deeply explore the model’s behavior change within each dimension caused by the appearance of other conflicting dimensions like our work does. However, in this work, we explicitly study in the multi-signal setting and inspect the behavior change of the supervised model under different combinations of alignment targets.
Appendix C The Complete Procedure for Performing Weak-to-Strong Preference Alignment
@@
Appendix D Concrete Mathematical Forms of All Weak-to-Strong Objectives
D.1 Reward Modeling Scenario
@@
(1) No Conflict: First, in order to explore the performance of the strong model it should have achieved without the conflicting alignment goal, we should obtain a weak-to-strong model trained under the weak supervision towards harmlessness only:
@@
D.2 Preference Alignment Scenario
@@
Appendix E Training Details
E.1 Code and Platform
@@
E.2 Training Details in The Reward Modeling Scenario
@@
E.3 Training Details in The Preference Alignment Scenario
@@
Appendix F Results of Test Accuracies in The Preference Alignment Scenario with SimPO
@@
Appendix G Weak-to-Strong Results on DPO
@@
Besides the main results on SimPO (Meng et al., 2024), we also conduct weak-to-strong preference alignment experiments on DPO (Rafailov et al., 2024). The detailed procedure is similar to SimPO and can be found in Appendix C. The major difference is the metric for determining the correctness of each prediction of the model and calculating the model’s confidence on each sample pair. In DPO, when using Eq. (9) to calculate the confidence of model $\bm{\theta}$ on $(x;y_{c},y_{r})$, the model logit $\pi_{\bm{\theta}}(y|x)$ on each completion $y$ is now normalized by a constant factor $L$ (instead of normalized by each response’s own sequence length) and is calculated as $\pi_{\bm{\theta}}(y|x)=\frac{1}{L}\sum\nolimits_{i=1}^{|y|}\log P_{\bm{\theta}}(y_{i}|x,y_{<i})$. This is because the original training objective of DPO is to directly enlarge the gap between the log sums of token probabilities over the chosen and rejected responses. Thus, the evaluation metric should be consistent to the training objective. We admit this may introduce a sequence length bias, but this is an inherent issue of DPO. However, the scale of original log sum of token probabilities is very huge, we need to divide it by a constant to ensure that the model’s confidence falls within a reasonable distribution. Choosing different constants for re-scaling may affect the confidence distribution, however, it is equivalent to selecting different confidence thresholds. The results in Appendix I validate that the deception pattern is independent of the choices of confidence thresholds. Here, we set $L$ to 50. The confidence threshold $T$ is fixed to 0.75. The results of generalization performance are displayed in Figure LABEL:fig:_accuracy_results_in_dpo, and the results of deception scores are shown in Figure 27. The results show that the weak-to-strong deception issue also exists in DPO setting.
Appendix H Experiments on LLaMA-3 and LLaMA-3.1
@@
Appendix I Deception Scores under Different Confidence Thresholds
In the main text, we display the patterns of deception scores in both two scenarios under the confidence threshold $T=0.75$. Choosing a different confidence threshold may affect the delineation of known and unknown areas of a target model. Here, we display the patterns of deception scores in both scenarios under three extra confidence thresholds ($T=0.70,0.80,0.85$). The results in the reward modeling and preference alignment (SimPO) scenarios are displayed in Figure LABEL:fig:_deception_scores_under_different_T_on_reward_modeling and Figure LABEL:fig:_deception_scores_under_different_T_in_preference_alignment, respectively. As we can see, though the concrete value of deception score in each experiment varies slightly under different $T$, the general pattern that the deception issue intensifies as the capability gap between the weak and strong models increases remains the same.
Appendix J Stronger Conflicting Strength Generally Leads to More Severe Deception
In main experiments, we fix the conflict strength factor $\alpha$ in the explicit conflict setting to 0.5. Here, we conduct extra experiments with a smaller $\alpha=0.25$, and put the comparison results in Figure 36. As we can expect, the deception score consistently increases as the degree of conflict increases.
Appendix K Explorations on The Spontaneous Weak-to-Strong Deception Issue in The No Conflict Setting
Figure 37: Absolute deception scores on the reward modeling task.
Figure 38: Absolute deception scores in the preference alignment scenario.
In the main text, we have studied the weak-to-strong deception issue in a realistic multi-objective alignment scenario. In this section, we make explorations on the spontaneous weak-to-strong deception issue: if current LLMs may spontaneously deceive weak supervisors even without being driven by conflicting targets. Specifically, we can compare the behavior change of the strong student in different knowledge areas when trained by no-conflict weak data with that trained by ground-truth data. Thus, we visualize the absolute deception score, which is calculated as the percentage of samples that are originally well-aligned under ground-truth supervision but now misaligned under weak supervision with no conflict, belonging to the Strong-Known and Weak-Unknown area. The full results in both reward modeling and preference alignment scenarios are in Figure 38 and Figure 38. We also put the results of absolute deception scores in the implicit conflict setting in same figures for comparison. Notice that in order to achieve fair comparisons, when calculating the absolute deception scores in the implicit conflict setting, the reference model used is also the ground truth strong model instead of the weak-to-strong model in the no conflict setting.
The main conclusions include: (1) The pattern of spontaneous weak-to-strong deception also exists, as most absolute deception scores are significantly larger than 0. (2) The spontaneous deception issue becomes more severe as the capability gap between weak and strong models increases. As we can see, supervised by the same weak data, the stronger model tends to make more mistakes in the Strong Known and Weak-Unknown area (refer to more visualizations in Appendix L). (3) The absolute deception scores in the no conflict setting are lower than deception scores under conflicting objectives, indicating that in realistic multi-objective alignment scenarios, the existence of conflicting optimization objectives exacerbates the deception issue. Considering the reason explained in Section 4.2 and the fact that in the real case where the alignment target is usually multi-objective, our main analysis under conflicting alignment targets aligns more closely with the real situation, but the analysis in the no conflict setting can be a good supplementary study.
Appendix L More Visualizations about The Dynamic Changes of Conflict Tax or Weak-to-Strong Tax
Here, we put the additional visualizations about the dynamic changes of conflict tax across all four knowledge areas when the weak model is GPT-2-XL in the explicit and implicit conflict settings in Figure LABEL:fig:_scatter_plots_in_explicit_conflict and Figure LABEL:fig:_scatter_plots_in_implicit_conflict, respectively.
Appendix M Full Results of High-Confidence Weak-to-Strong Preference Alignment
@@ Appendix I Deception Scores under Different Confidence Thresholds
Appendix N Details in Bootstrapping Experiments
@@
Appendix O Preliminary Experiments on Honest Alignment
Besides the harmlessness goal considered in the main experiments, here, we perform preliminary experiments by regarding honesty as the target alignment dimension, and explore potential weak-to-strong deception issue when the conflicting target helpfulness appears. The motivation is, honesty requires the model to refuse the questions it does not know while helpfulness requires the model to provide helpful information on any user question. We select and filter the honesty data from UnknownBench (Liu et al., 2023), where the prompts are unanswerable questions, preferred responses are from gpt-4-0613, dispreferred responses are from Llama-2-13B-Chat. After filtering, we finally obtain 400 samples for ground truth training data, 400 samples for weak data and 400 samples for testing data. We perform experiments on GPT-2-series models in the preference alignment scenario with SimPO in the implicit conflict setting, where we include 2,000 helpful samples for the conflicting objective. The batch size for training is 16, while other experimental settings are kept as the same as that in Section 5.2. The results are in Figure LABEL:fig:_results_on_honesty. The weak-to-strong deception issue also exists in this honest alignment setting.
Appendix P The Effect of Adaptive Supervision Method on Mitigating Deception
In the main text, we discuss the potential deception mitigation strategy by only using correct and high-confidence samples for weak-to-strong alignment, but obtain the negative results. Here, we conduct extra experiments on an adaptive supervision method. The motivation is to dynamically down-weights the importance of low-confidence samples predicted by the weak model. Thus, we design an adaptive loss function as
$\tilde{\bm{\theta}}_{s}^{w}=$ $\mathop{\arg\min}_{\bm{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\left(\mathcal{L}_{CE}\big{(}M_{\bm{\theta}_{s}}(x),M_{\bm{\theta}_{w}^{gt}}(x)\big{)}\cdot|2M_{\bm{\theta}_{w}^{gt}}(x)-1|\right),$ (16)
where $|2M_{\bm{\theta}_{w}^{gt}}(x)-1|$ is a re-weighting factor that relatively down-weights low-confidence samples.
Figure 48: The comparison results of deception scores between naive weak-to-strong loss and adaptive weak-to-strong loss.
We conduct experiments in the reward modeling task under implicit conflict setting, as strong models are not expected to get the probabilities in the preference alignment scenario. Other experimental settings are kept as the same as that in Section 4.1. The comparison results between this adaptive loss and naive weak-to-strong loss in Eq. (4) are put in Figure 48. As we can see, the deception mitigation performance of adaptive loss is also limited.
Appendix Q Case Study on Weak-to-Strong Deception